# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 3. Modular Forms on $\textup{SO}(2,n)$
We will now define a special class of Fourier–Jacobi forms, called Poincaré series. To begin with, we define the support of the lattice $L=\mathbb{Z}^{n}$ to be
@@
$\textup{supp}(L,\phi):=\left\{(D,r)\mid D\in\mathbb{Q}_{\leq 0},r\in L^{*},D\equiv\frac{1}{2}S[r]\pmod{\mathbb{Z}}\right\}$.
Now, if we write $r\longmapsto S^{-1}r$ with $r\in L$, we get
$qD\equiv\frac{1}{2}qS^{-1}[r]\pmod{q\mathbb{Z}}$,
@@
where $q$ is the level of the lattice. This then implies $D\in\frac{1}{q}\mathbb{Z}$. Hence, by writing $D\longmapsto D/q$, we get that equivalently
@@
$\textup{supp}(L,\phi)=\left\{(D/q,S^{-1}r)\mid D\in\mathbb{Z}_{\leq 0},r\in L,D\equiv\frac{1}{2}qS^{-1}[r]\pmod{q}\right\}$.
@@
$\widetilde{\textup{supp}}(L,\phi):=\{(D,r)\in\mathbb{Z}_{\leq 0}\times L\mid(D/q,S^{-1}r)\in\textup{supp}(L,\sigma)\}.$ (9)
and in the following this is the set we will use.
@@
If now $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$, we define the Poincaré series $P_{k,D,r}$ as in [Moc19, Definition 2.2]. If $k>n/2+2$, $P_{k,D,r}$ is absolutely and uniformly convergent on compact subsets of $\mathbb{H}\times(L\otimes\mathbb{C})$ ($\mathbb{H}$ the usual upper half plane) and defines an element of $S_{k}(L,\phi)$ (see [Moc19, Theorem 2.3, (i)]).
Finally, if $N\geq 1$, we denote by $V_{N}:S_{k}(L,\phi)\longrightarrow S_{k}(L,N\phi)$ the index-raising operator defined in [Sug95, Lemma 6.1] and [Moc19, Definition 4.25]. In particular, for $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$, we define
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{k,D,r})(\tau,z)e(N\tau^{\prime}).$ (10)

### v2

@@ 5. Poincaré Series
We will now define a special class of Fourier–Jacobi forms, called Poincaré series, associated to the lattices $L_{\mathfrak{m}}$ above. To begin with, the support of the lattice $L_{\mathfrak{m}}$ in $V$ is
@@
$\textup{supp}(L_{\mathfrak{m}},\phi):=\left\{(D_{\mathfrak{m}},r_{\mathfrak{m}})\mid D_{\mathfrak{m}}\in\mathbb{Q}_{\leq 0},r_{\mathfrak{m}}\in L_{\mathfrak{m}}^{*},D_{\mathfrak{m}}\equiv\frac{1}{2}S[r_{\mathfrak{m}}]\pmod{\mathbb{Z}}\right\}$.
Now, since $L_{\mathfrak{m}}^{*}=S^{-1}A_{\mathfrak{m}}^{-t}\mathbb{Z}^{2}$, we write $r_{\mathfrak{m}}\longmapsto S^{-1}A_{\mathfrak{m}}^{-t}r_{\mathfrak{m}}$, with $r_{\mathfrak{m}}\in\mathbb{Z}^{2}$. Hence, the congruence becomes
$q_{\mathfrak{m}}D_{\mathfrak{m}}\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[r_{\mathfrak{m}}]\pmod{q_{\mathfrak{m}}\mathbb{Z}}$,
@@
where $q_{\mathfrak{m}}$ is the level of the lattice $L_{\mathfrak{m}}$. This then implies $D_{\mathfrak{m}}\in\frac{1}{q_{\mathfrak{m}}}\mathbb{Z}$. Hence, by writing $D_{\mathfrak{m}}\longmapsto D_{\mathfrak{m}}/q_{\mathfrak{m}}$, we get that equivalently
@@
$\textup{supp}(L,\phi)=\left\{(D_{\mathfrak{m}}/q_{\mathfrak{m}},S^{-1}A_{\mathfrak{m}}^{-t}r_{\mathfrak{m}})\mid D_{\mathfrak{m}}\in\mathbb{Z}_{\leq 0},r_{\mathfrak{m}}\in\mathbb{Z}^{2},D_{\mathfrak{m}}\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[r_{\mathfrak{m}}]\pmod{q_{\mathfrak{m}}}\right\}$.
@@
(5.1) $\widetilde{\textup{supp}}(L_{\mathfrak{m}},\phi):=\{(D_{\mathfrak{m}},r_{\mathfrak{m}})\in\mathbb{Z}_{\leq 0}\times\mathbb{Z}^{2}\mid(D_{\mathfrak{m}}/q_{\mathfrak{m}},S^{-1}A_{\mathfrak{m}}^{-t}r_{\mathfrak{m}})\in\textup{supp}(L_{\mathfrak{m}},\phi)\}.$
and in the following, this is the set we will use.
@@
If now $(D_{\mathfrak{m}},r_{\mathfrak{m}})\in\widetilde{\textup{supp}}(L_{\mathfrak{m}},\phi)$, we define the Poincaré series $P(k,D_{\mathfrak{m}},r_{\mathfrak{m}})$ as in [jacobi_lattice, Definition 2.2]. If $k>n/2+2$, $P(k,D_{\mathfrak{m}},r_{\mathfrak{m}})$ is absolutely and uniformly convergent on compact subsets of $\mathbb{H}\times(L_{\mathfrak{m}}\otimes\mathbb{C})$ ($\mathbb{H}$ the usual upper half plane) and defines an element of $S_{k}(L_{\mathfrak{m}},\phi)$ (see [jacobi_lattice, Theorem 2.3, (i)]).
For $N\geq 1$, we denote by $V_{N}:S_{k}(L_{\mathfrak{m}},\phi)\longrightarrow S_{k}(L_{\mathfrak{m}},N\phi)$ the index-raising operator defined in
[jacobi_lattice, Definition 4.25]. In particular, for $(D_{\mathfrak{m}},r_{\mathfrak{m}})\in\widetilde{\textup{supp}}(L_{\mathfrak{m}},\phi)$, we define
(5.2) $\mathcal{P}_{\mathfrak{m}}(\tau^{\prime},z,\tau):=\mathcal{P}(k,D_{\mathfrak{m}},r_{\mathfrak{m}})(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P(k,D_{\mathfrak{m}},r_{\mathfrak{m}}))(\tau,z)e(N\tau^{\prime}).$